OpenStudy (anonymous):

solve the equation 1/2x+4=-2/3x+1/2

OpenStudy (anonymous):

1/2x + 4 = -2/3x + 1/2 First, multiply everything by 3, to get rid of the fraction -2/3x: 3/2x + 12 = -2x + 3/2 Then, multiply everything by 2, to get rid of all remaining fractions: 3x + 24 = -4x + 3 Then, add 4x to both sides, to get rid of -4x: 7x + 24 = 3 Subtract 24 from both sides: 7x = -21 Divide by 7: x = -3 Hope that helped!

OpenStudy (anonymous):

thank you very Much!